In the combustion of hydrogen gas, hydrogen reacts with oxygen from the air to form water vapor.

hydrogen+oxygen⟶water

If you burn 47.4 g of hydrogen and produce 423 g of water, how much oxygen reacted?
Answer #1

This can be explained based on two methods,

1)

Combustion reaction of hydrogen gas with oxygen can be written as follows. Two moles of Hydrogen react with one mole of Oxygen to form two moles of . That means 4 grams of hydrogen react with 32 grams of Oxygen to result 36 grams of water.

379.2 g of oxygen will be required for the combustion of 47.4 gram of hydrogen which will result 426.6 g of water ( not exactly 423 g).

2) Law of conservation of mass.

If 47.4g of hydrogen react with oxygen to form exactly 423 g of water, law of conservation of mass cane be used to find the amount of water reacted. According to this "Total mass of reactants will be equal to the total mass of products".

Total mass of reactants= total mass of H​​​​​​2 + total mass of O​​​2 =47.4+X , where X= mass of Oxygen

Total mass of products= mass of water = 423 g

From these equations,

Mass of Oxygen= Mass of water - mass of hydrogen= 423-47.4=375.6 g.

375.6 g of Oxygen react with 47.4 g of hydrogen to form 423 g of water .